You are given the sample mean and the population standard deviation. Use this information to construct the​ 90% and​ 95% confidence intervals for the population mean. Interpret the results and compare the widths of the confidence intervals. If​ convenient, use technology to construct the confidence intervals. A random sample of 60 home theater systems has a mean price of ​$113.00. Assume the population standard deviation is ​$17.90. Construct a​ 90% confidence interval for the population mean
Answer #1

Solution :

Given that,

(a)

At 90%

Z\alpha/2 = 1.645

Margin of error = E = Z\alpha/2* (\sigma /\sqrtn)

= 1.645 * (17.90 / \sqrt 60)

= 3.80

At 90% confidence interval estimate of the population mean is,

\bar x - E < \mu < \bar x + E

113 - 3.80 < \mu < 113 + 3.80

109.2 < \mu < 116.80

(109.2  , 116.80)

(ii)

At 95%

Z\alpha/2 = 1.96

Margin of error = E = Z\alpha/2* (\sigma /\sqrtn)

= 1.96 * (17.90 / \sqrt 60)

=4.53

At 95% confidence interval estimate of the population mean is,

\bar x - E < \mu < \bar x + E

113 - 4.53 < \mu < 113 + 4.53

108.47 < \mu < 117.53

(109.2  , 116.80)